# Random expansions of trees with bounded height

### v1

@@ 1. Introduction
When building a model for probabilistic inference we may not only want to take into account properties and relations which are uncertain and described by (conditional) probabilities, but also “background information” that is certain. A common kind of background information is information about a hierarchical structure, and such information can be described by a tree. As an example of a tree that represents a hierarchical structure, consider a tree in which the children of the root represent some communities, where the children of a community represent subcommunities, and the children of a subcommunity represent individuals. The height of this hierarchy, or tree, is 3 (or 4, depending on our conventions). But the number of individuals in even the smallest subcommunity that we consider may be much larger than 3. This article considers hierarchical background information represented by a tree such that the height of the tree is very much smaller than the minimal number of children of a nonleaf of the tree.
@@
On top of the background information, represented by the example tree described above, we may consider information that is given by (conditional) probabilities and is formally represented by (parametrized) 0/1-valued (i.e. false/true-valued) random variables. So we may have random variables (also viewed as atomic logical formulas) $P_{1}(x)$, $P_{2}(x)$, and $P_{3}(x)$, representing some property that a community, subcommunity, respectively, individual, may have. The probability that a subcommunity has the property $P_{2}$ may (for example) depend on whether the community which it is part of has property $P_{1}$. Or alternatively, the probability that a community has $P_{1}$ may depend on the proportion of its subcommunities that have $P_{2}$ (and/or the proportion of individuals in the community that have $P_{3}$). We may also have a random variable $R(x,y)$, representing some relationship between objects, for example the perception (good/bad) that $x$ has of $y$. Now, for an individual $x$, the probability that $R(x,y)$ holds may depend on properties of the (sub)community that $x$ belongs to and on properties of $y$ and/or its (sub)community; it may also depend on whether $x$ and $y$ belong to the same (sub)community.
@@
$PLA^{*}$ is a logic such that the (truth) values of its formulas can be any number in the unit interval $[0,1]$, so it is suitable for expressing probabilities. $PLA^{*}$ achieves this partly by considering every $c\in[0,1]$ as an atomic formula, and partly by using aggregation functions which take finite sequences of reals in $[0,1]$ as input and outputs a real number in $[0,1]$. An example of an aggregation function is the arithmetic mean, or average, which returns the average of a finite sequence of reals in the unit interval. So for example, if $R$ is a unary relation symbol then there is a $PLA^{*}$-formula (without free variables) such that its value is the proportion of elements in the domain which satisfy $R(x)$. Since the aggregation functions maximum and minimum are available there is a $PLA^{*}$ formula which expresses that at least one (alternatively all) element in the domain satisfies $R(x)$. By using the aggregation functions maximum and minimum one can show that every query on finite structures which can be expressed by a first-order formula can be expressed by a $PLA^{*}$-formula (Lemma 3.13). Hence, $PLA^{*}$ subsumes first-order logic. For a formula $\varphi(x_{1},\ldots,x_{k})$ of $PLA^{*}$, finite structure $\mathcal{A}$ (with matching signature/vocabulary), and $a_{1},\ldots,a_{k}\in A$, we let $\mathcal{A}(\varphi(a_{1},\ldots,a_{k}))$ denote the value of $\varphi(a_{1},\ldots,a_{k})$ in $\mathcal{A}$.
@@
The precise way that we will mathematically represent the kind of probabilistic (or uncertain) information described above together with some background information, in the form of a tree, is as follows. We let $\Delta$ be a fixed positive integer and we let $\mathcal{T}_{n}$, where $n$ is a positive integer, be a tree with maximum height $\Delta$ that represents the hierarchical background information. Since we are interested in what happens for (arbitrarily) large trees where nonleaves have many children, we assume that, for every positive integer $k$, if $n$ is large enough then every nonleaf in $\mathcal{T}_{n}$ has at least $k$ children. Each $\mathcal{T}_{n}$ can be represented by a first-order $\tau$-structure where $\tau=\{E\}$ and $E$ is a binary relation symbol of arity 2. Then let $\sigma$ be a finite relational signature that includes $\tau$. In the example above we have $\sigma=\{E,P_{1},P_{2},P_{3},R\}$. Then let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, where we think of each such expansion $\mathcal{A}$ as a “possible world”, and let $\mathbb{P}_{n}$ be the probability distribution on $\mathbf{W}_{n}$ induced by a $PLA^{*}(\sigma)$-network $\mathbb{G}$ as defined in Definition 6.9 .
@@
So given a sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ of trees (of the described kind) and a $PLA^{*}(\sigma)$-network $\mathbb{G}$, we have a precise mathematical setting of computing the probability of a query (event) on $\mathbf{W}_{n}$ that can be expressed by a $PLA^{*}$-formula. The probability of a query depends on both on the hierarchical information given by $\mathcal{T}_{n}$ and by $\mathbb{G}$. However, if we use the brute force way of computing the probability of a query, namely that we compute the probability of every structure in $\mathbf{W}_{n}$ in which the query is true and then add all such probabilities, then the time needed for the computation will be exponential in the number of vertices of the tree $\mathcal{T}_{n}$. In other words, the brute force method of computing the probability of a query does not scale well to large trees $\mathcal{T}_{n}$. This motivates the search for other methods of computing, or estimating, the probability of a query in the described context.
@@
    certain conditions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ (the stronger Assumption 6.3 and the milder Assumption 8.3),
@@
The main results, stated in Section 8.1, have the general form described above. We first prove Theorem 8.1. Then we make some observations which allow us to easily obtain some corollaries which are variations of Theorem 8.1. The essence of these corollaries is (i) that if we consider stronger assumptions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ (i.e. less general probability distributions), then we get results that apply to more queries, and (ii) if we loosen the assumptions on the sequence of trees $\mathbf{T}$ then we get results that apply to less, but still interesting, queries.
@@
Related work. Instead of bounding the height of the trees, as we do in this article, one could consider trees with a bound on the number of children that vertices may have. Such sequences of trees are covered by the (more general) context considered by Koponen in [20], where probability distributions where defined by $PLA^{*}(\sigma)$-networks and queries by $PLA^{*}$-formulas. Other work that we are aware of on logical convergence laws, and related issues, in the setting of expansions of “base structures” uses first-order logic (or $L_{\infty,\omega}^{\omega}$) as the query language. Baldwin [4] and Shelah [27] considered a base sequence where each $\mathcal{B}_{n}$ is a directed path of length $n$. In both cases $\mathbf{W}_{n}$ consists of all expansions of $\mathcal{B}_{n}$ to a new binary relation symbol $R$ interpreted as an irreflexive and symmetric relation. In [4] the probability distribution on $\mathbf{W}_{n}$ is defined by letting an $R(x,y)$ be true with probability $n^{-\alpha}$ for some irrational $\alpha\in(0,1)$, independently of whether it is true for other pairs. In [27] the probability of $R(x,y)$ is $d^{-\alpha}$ where $d$ is the distance in $\mathcal{B}_{n}$ between $x$ and $y$ and $\alpha\in(0,1)$ is irrational. Lynch [25, Corollary 2.16] and later Abu Zaid, Dawar, Grädel and Pakusa [1] and Dawar, Grädel and Hoelzel [10] considered the context where $\mathcal{B}_{n}$ is a product of finite cyclic groups and the probability distribution considered is the uniform one. Ahlman and Koponen [2] considered base structures $\mathcal{B}_{n}$ with a (nicely behaved) pregeometry and a probability distribution defined by a kind of stochastic block model where the “blocks” correspond to subspaces. Lynch [25] formulated a condition ($k$-extendibility) that makes sense for base sequences $\mathbf{B}$ in general and which guarantees that a convergence law holds for first-order logic and the uniform probability distribution.
@@
Besides this article and [20], other studies that use a a PGM for generating a probability distribution and that prove some result about asymptic probabilities do not consider background information in the form of a sequence of nontrivial base structures. But in the context of no background information, when we let $\sigma$ be a finite and relational signature, $\mathbf{W}_{n}$ the set of all $\sigma$-structures with domain $\{1,\ldots,n\}$, $\mathbb{P}_{n}$ a probability distribution induced by a PGM, there are some studies. The most closely related to this one are [21, 22, 23] by Koponen and Weitkämper and [14] by Jaeger. But we also have [19] by Koponen, [28, 29] by Weitkämper, and [9] by Cozman and Maua. The first result about “asymptotic equivalence” between complex first-order formulas and simpler formulas was probably given by Glebskii et. al. [13] and was used to prove the classical zero-one law for first-order logic, also proved independently by somewhat different methods by Fagin [12]. Since the pioneering work in [13] and [12] a large number of studies on logical convergence laws and related issues have been conducted in the fields of finite model theory and probabilistic combinatorics, but those studes consider less flexible ways of generating probability distributions than PGMs (most often the uniform distribution).
@@
Organization. Section 2 specifies some notation and terminology that will be used, as well as a couple of probability theoretic results. Section 3 defines the formal logic $PLA^{*}$, and some related notions, that will be used in PGMs and for defining queries (events). Section 4 describes a general method (introduced in [23]) for “asymptotically eliminating aggregation functions”. Section 5 gives the necessary background about directed acyclic graphs and trees, and some other notions related to trees, that will be used in the rest of the article. Section 6 specifies the assumptions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ that we make and also explains why we make these assumptions. It then defines the notion of a $PLA^{*}$-network, which is the kind of PGM that we use, and how it induces a probability distribution on the set $\mathbf{W}_{n}$ of all expansions to $\sigma$ of $\mathcal{T}_{n}$. In section 7 we prove technical results that imply that the assumptions of Theorem 4.8 in Section 4 are satisfied. Therefore we can, in the last subsection of Section 7, use Theorem 4.8 and the results from earlier subsections of Section 7 to prove a result about “asymptotic elimination of aggregation functions”. In Section 8 we use induction on the height of the $PLA^{*}$-network and the results from Section 7 to prove our main results.
@@ 2. Preliminaries
Suppose that $\mathbb{P}$ is a probability measure on a set $\Omega$. Let $X\subseteq\Omega$ and $Y\subseteq\Omega$ be measurable. Also suppose that $Y=Y_{1}\cup\ldots\cup Y_{k}$, $Y_{i}\cap Y_{j}$ if $i\neq j$, and that each $Y_{i}$ is measurable. If $\alpha\in[0,1]$, $\varepsilon>0$, and $\mathbb{P}(X\ |\ Y_{i})\in[\alpha-\varepsilon,\alpha+\varepsilon]$ for all $i=1,\ldots,k$, then $\mathbb{P}(X\ |\ Y)\in[\alpha-\varepsilon,\alpha+\varepsilon]$.

### v2

@@ 1. Introduction
Logical convergence laws have been proved or disproved in various contexts since the pioneering work of Glebskii et al [13] and, independently, Fagin [12]. Since finite structures in the sense of first-order logic can represent, for example, relational databases, such convergence laws may have impact outside of mathematics itself. When a convergence law applies to a sentence of a formal logic one can use random sampling of structures on a sufficiently large domain to get an estimate of the probability of the sentence on all large enough domains and the estimate will, with high probability, be as close to the actual probability as we like (according to some predetermined error marginal). Moreover, sometimes a logical convergence law can be proved (as in [13]) by showing that quantifiers/aggregations can be removed from formulas step by step to produce a new simpler (e.g. quantifier-free) formula which is “asymptotically equivalent” to the original one, and the simpler formula can be evaluated in time which is independent of the domain size. In theory this gives a deterministic algorithm for estminating the probability of a sentence (or formula with some parameters/constants) which is independent of the domain size. In practice, properties of the (generalized) quantifiers or aggregation functions that are involved in the elimination will of course influence the computational complexity.
Most studies of logical convergence laws have studied contexts in which all relations are modelled probabilistically; intuitively speaking they are “uncertain”. When building a model for inference we may not only want to take into account properties and relations which are uncertain, but also “background information” that is certain. Background information can have many different forms, but arguably some forms of background information are more common, or more important, than other forms. Information that categorizes objects into classes, subclasses and so on is prevalent both in theory and in the physical world. A related form of information is that of a hierarchical structure among objects. Both forms of information can be represented by trees: the root represents the class of all objects (or the top of the hierarchy), the children of the root the classes of the first subdivision into subclasses (or the next level of the hierarchy), and so on. This is a motivation to consider logical convergence laws for random structures that expand a tree, that is, the tree is fixed but in addition we have random relations and the probability that such a relation, say $R(a,b)$, holds may depend on the positions of $a$ and $b$ in the tree.
@@
As an example, consider a tree in which the children of the root represent some communities, where the children of a community represent subcommunities, and the children of a subcommunity represent individuals. On top of the background information, represented by the tree, we may consider information that is given by (conditional) probabilities and is formally represented by (parametrized) 0/1-valued (i.e. false/true-valued) random variables. So we may have random variables (also viewed as atomic logical formulas) $P_{1}(x)$, $P_{2}(x)$, and $P_{3}(x)$, representing some property that a community, subcommunity, respectively, individual, may have. The probability that a subcommunity has the property $P_{2}$ may (for example) depend on whether the community which it is part of has property $P_{1}$. Or alternatively, the probability that a community has $P_{1}$ may depend on the proportion of its subcommunities that have $P_{2}$ (and/or the proportion of individuals in the community that have $P_{3}$). We may also have a random variable $R(x,y)$, representing some relationship between objects, for example the perception (good/bad) that $x$ has of $y$. Now, for an individual $x$, the probability that $R(x,y)$ holds may depend on properties of the (sub)community that $x$ belongs to and on properties of $y$ and/or its (sub)community; it may also depend on whether $x$ and $y$ belong to the same (sub)community.
@@
$PLA^{*}$ is a logic such that the (truth) values of its formulas can be any number in the unit interval $[0,1]$, so it is suitable for expressing probabilities. $PLA^{*}$ achieves this partly by considering every $c\in[0,1]$ as an atomic formula, and partly by using aggregation functions which take finite sequences of reals in $[0,1]$ as input and outputs a real number in $[0,1]$. An example of an aggregation function is the arithmetic mean, or average, which returns the average of a finite sequence of reals in the unit interval. So for example, if $R$ is a unary relation symbol then there is a $PLA^{*}$-formula (without free variables) such that its value is the proportion of elements in the domain which satisfy $R(x)$. By using the aggregation functions maximum and minimum one can show that every query on finite structures which can be expressed by a first-order formula can be expressed by a $PLA^{*}$-formula (Lemma 3.13). Hence, $PLA^{*}$ subsumes first-order logic. For a formula $\varphi(x_{1},\ldots,x_{k})$ of $PLA^{*}$, finite structure $\mathcal{A}$ (with matching signature/vocabulary), and $a_{1},\ldots,a_{k}\in A$, we let $\mathcal{A}(\varphi(a_{1},\ldots,a_{k}))$ denote the value of $\varphi(a_{1},\ldots,a_{k})$ in $\mathcal{A}$.
@@
Since we are interested in large trees we let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be a sequence of trees where the number of vertices of $\mathcal{T}_{n}$ tends to infinity as $n\to\infty$. Each $\mathcal{T}_{n}$ can be represented by a first-order $\tau$-structure where $\tau=\{E\}$ and $E$ is a binary relation symbol of arity 2. Let $\sigma$ be a finite relational signature that includes $\tau$. In the example above we have $\sigma=\{E,P_{1},P_{2},P_{3},R\}$. Then let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, where we think of each such expansion $\mathcal{A}$ as a “possible world”, and let $\mathbb{P}_{n}$ be the probability distribution on $\mathbf{W}_{n}$ induced by a $PLA^{*}(\sigma)$-network $\mathbb{G}$ as defined in Definition 6.9.
@@
Now we have a precise mathematical setting in which we can speak about the probability of a query (event) on $\mathbf{W}_{n}$ that can be expressed by a $PLA^{*}$-formula. The probability of a query depends (in general) on both the hierarchical information given by $\mathcal{T}_{n}$ and by $\mathbb{G}$. If we use the brute force way of computing the probability of a query, namely that we compute the probability of every structure in $\mathbf{W}_{n}$ in which the query is true and then add all such probabilities, then the time needed for the computation will be exponential in the number of vertices of the tree $\mathcal{T}_{n}$. In other words, the brute force method of computing the probability of a query does not scale well to large trees $\mathcal{T}_{n}$. This motivates the search for other methods of computing, or estimating, the probability of a query in the described context. We do not expect to get a general convergence (or nonconvergence) result that covers every kind of sequence $\mathbf{T}$, so additional assumptions on $\mathbf{T}$ will be imposed. Koponen’s results in [20] apply to sequences $\mathbf{T}$ where for some $\Delta$ and all $n$ each vertex in $\mathcal{T}_{n}$ has at most $\Delta$ children. In this work we instead assume that for some $\Delta$ and all $n$ the height of $\mathcal{T}_{n}$ is at most $\Delta$. It will also be assumed that for every $k$ and all large enough $n$, every nonleaf in $\mathcal{T}_{n}$ has at least $k$ children. Such trees can (for large $n$) describe hierarchies (categorizations) in with few levels compared to the number of objects of each type in the hierarchy (or few partitions into subcategories compared to the number of objects in each category). In the above example the height of the tree, is 3 (or 4, depending on our conventions), but the number of individuals in even the smallest subcommunity that we consider may be much larger than 3.
@@
    certain additional conditions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ (the stronger Assumption 6.3 and the milder Assumption 8.3),
@@
Although condition (1) in the form of Assumption 6.3 has a rather technical statement, the general intuition is just that each tree $\mathcal{T}_{n}$ is “sufficiently homogeneous” in the sense that for every nonroot vertex $a$ and subtree $\mathcal{T}^{\prime}$ of $\mathcal{T}_{n}$ rooted in $a$ there are “sufficiently many” siblings $b$ of $a$ such that the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$ is rougly the same as the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$. For example, suppose that, in a clinical test say, a group of $n$ persons, where $n$ is large, is first divided into “many” groups of roughly the same (large) size, then each group is subdivided into roughly equally many subclasses of roughly the same (still large) size, and so on for, say 5, subdivisions. (The divisions may be based on putting people with similar values with respect to various measurements into the same group.) Given that $n$, “many”, and “large” are large enough, the tree $\mathcal{T}_{n}$ representing the subdivisions will be of the kind considered in (1). More precisely (according to Assumption 6.3), it suffices that “many” and “large” means “growing faster” than $c\ln n$ for every positive contant $c$.
@@
The main results, stated in Section 8.1, have the general form described above. We first prove Theorem 8.1. Then we make some observations which allow us to easily obtain some corollaries which are variations of Theorem 8.1. The essence of these corollaries is (i) that if we consider stronger assumptions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ (i.e. less general probability distributions), then we get results that apply to more queries, and (ii) if we loosen the assumptions on the sequence of trees $\mathbf{T}$ then we get results that apply to fewer, but still interesting, queries.
@@
Related work. Instead of bounding the height of the trees, as we do in this article, one could consider trees with a bound on the number of children that vertices may have. Such sequences of trees are covered by the (more general) context considered by Koponen in [20], where probability distributions were defined by $PLA^{*}(\sigma)$-networks and queries by $PLA^{*}$-formulas. Other work that we are aware of on logical convergence laws, and related issues, in the setting of expansions of “base structures” uses first-order logic (or $L_{\infty,\omega}^{\omega}$) as the query language. Baldwin [4] and Shelah [27] considered a base sequence where each $\mathcal{B}_{n}$ is a directed path of length $n$. In both cases $\mathbf{W}_{n}$ consists of all expansions of $\mathcal{B}_{n}$ to a new binary relation symbol $R$ interpreted as an irreflexive and symmetric relation. In [4] the probability distribution on $\mathbf{W}_{n}$ is defined by letting an $R(x,y)$ be true with probability $n^{-\alpha}$ for some irrational $\alpha\in(0,1)$, independently of whether it is true for other pairs. In [27] the probability of $R(x,y)$ is $d^{-\alpha}$ where $d$ is the distance in $\mathcal{B}_{n}$ between $x$ and $y$ and $\alpha\in(0,1)$ is irrational. Lynch [25, Corollary 2.16] and later Abu Zaid, Dawar, Grädel and Pakusa [1] and Dawar, Grädel and Hoelzel [10] considered the context where $\mathcal{B}_{n}$ is a product of finite cyclic groups and the probability distribution considered is the uniform one. Ahlman and Koponen [2] considered base structures $\mathcal{B}_{n}$ with a (nicely behaved) pregeometry and a probability distribution defined by a kind of stochastic block model where the “blocks” correspond to subspaces. Lynch [25] formulated a condition ($k$-extendibility) that makes sense for base sequences $\mathbf{B}$ in general and which guarantees that a convergence law holds for first-order logic and the uniform probability distribution.
@@
Besides this article and [20], other studies that use a PGM for generating a probability distribution and that prove some result about asymptotic probabilities do not consider background information in the form of a sequence of nontrivial base structures. But in the context of no background information, when we let $\sigma$ be a finite and relational signature, $\mathbf{W}_{n}$ the set of all $\sigma$-structures with domain $\{1,\ldots,n\}$, $\mathbb{P}_{n}$ a probability distribution induced by a PGM, there are some studies. The most closely related to this one are [21, 22, 23] by Koponen and Weitkämper and [14] by Jaeger. But we also have [19] by Koponen, [28, 29] by Weitkämper, and [9] by Cozman and Maua. The first result about “asymptotic equivalence” between complex first-order formulas and simpler formulas was probably given by Glebskii et. al. [13] and was used to prove the classical zero-one law for first-order logic, also proved independently by somewhat different methods by Fagin [12]. Since the pioneering work in [13] and [12] a large number of studies on logical convergence laws and related issues have been conducted in the fields of finite model theory and probabilistic combinatorics, but those studies consider less flexible ways of generating probability distributions than PGMs (most often the uniform distribution).
The present work and [21, 22, 20] consider similar formalisms for inducing probability and for defining queries/events, so we point out some (unavoidable) differences. In [21, 22] all relations are uncertain (probabilistically modelled) and every formula to which the main results(s) applies is asymptotically equivalent to a formula without aggregation functions and without quantifiers. In the present work and in [20] which also considers deterministic background information, represented by “base structures”, it is only shown that every formula to which some of the main results applies is asymptotically equivalent to a formula which may use (only) “local” aggregations/quantifications where the notion of “locality” is adapted to the kind of base structures considered, which in the present work are trees. It is not possible to remove the use of local aggregations/quantifications in this work (or in [20]) because aggregation/quantifier-free formulas cannot describe the structure of trees (or the base structures considered in [20]). Besides this difference compared with [21, 22], another difference is that in the present work (and [20]) we only asymptotically eliminate aggregations that range over elements, or tuples of elements, that satisfy some constraints with respect to the underlying base structure. It seems that one can not in general remove this assumption because probabilities may depend on the base structure.
@@
Organization. Section 2 specifies some notation and terminology that will be used, as well as a couple of probability theoretic results. Section 3 defines the formal logic $PLA^{*}$, and some related notions, that will be used in PGMs and for defining queries (events). Section 4 describes a general method (introduced in [23]) for “asymptotically eliminating aggregation functions”. Section 5 gives the necessary background about directed acyclic graphs and trees, and some other notions related to trees, that will be used in the rest of the article. Section 6 specifies the assumptions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ that we make and also explains why we make these assumptions. It then defines the notion of a $PLA^{*}$-network, which is the kind of PGM that we use, and how it induces a probability distribution on the set $\mathbf{W}_{n}$ of all expansions to $\sigma$ of $\mathcal{T}_{n}$. In section 7 we prove technical results that imply that the assumptions of Theorem 4.8 in Section 4 are satisfied. Therefore, we can, in the last subsection of Section 7, use Theorem 4.8 and the results from earlier subsections of Section 7 to prove a result about “asymptotic elimination of aggregation functions”. In Section 8 we use induction on the height of the $PLA^{*}$-network and the results from Section 7 to prove our main results.
@@ 2. Preliminaries
Suppose that $\mathbb{P}$ is a probability measure on a set $\Omega$. Let $X\subseteq\Omega$ and $Y\subseteq\Omega$ be measurable. Also suppose that $Y=Y_{1}\cup\ldots\cup Y_{k}$, $Y_{i}\cap Y_{j}=\emptyset$ if $i\neq j$, and that each $Y_{i}$ is measurable. If $\alpha\in[0,1]$, $\varepsilon>0$, and $\mathbb{P}(X\ |\ Y_{i})\in[\alpha-\varepsilon,\alpha+\varepsilon]$ for all $i=1,\ldots,k$, then $\mathbb{P}(X\ |\ Y)\in[\alpha-\varepsilon,\alpha+\varepsilon]$.